# Revealing Language Model Trajectories via Kullback-Leibler Divergence

### v1

@@ 33 DeepMount00/Llama-3-8b-Ita meta-llama/Meta-Llama-3-8B llama-3 2 179401
37 openchat/openchat-3.6-8b-20240522 Wang et al. (2024a) meta-llama/Meta-Llama-3-8B llama-3 2 10464
38 rinna/llama-3-youko-8b Andonian et al. (2021); Mitsuda et al. (2024); AI@Meta (2024); Sawada et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 1468
@@
40 lightblue/suzume-llama-3-8B-multilingual Devine (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 16442
@@
42 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct AI@Waktaverse (2024); AI@Meta (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2305
43 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 55718
@@ 44 SJ-Donald/llama3-passthrough-chat meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2241
45 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA Basile et al. (2023); AI@Meta (2024); Polignano et al. (2024) meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 5995
46 NousResearch/Hermes-2-Theta-Llama-3-8B "Teknium" et al. (2024a) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 12392
47 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B Gallego (2024) NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 10273
@@
52 cypienai/cymist-2-v02-SFT Lacoste et al. (2019) mistralai/Mistral-7B-v0.1 mistral 2 2717
@@
54 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 12911
55 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 294019
56 ibm/merlinite-7b Sudalairaj et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 11156
@@
58 Intel/neural-chat-7b-v3-1 Mukherjee et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 3976
59 kaist-ai/mistral-orpo-capybara-7k Hong et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 4821
@@
61 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistralai/Mistral-7B-v0.1 mistral 2 1370245
62 NousResearch/Hermes-2-Pro-Mistral-7B "interstellarninja" et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 14586
63 NousResearch/Nous-Hermes-2-Mistral-7B-DPO "Teknium" et al. (2024c) mistralai/Mistral-7B-v0.1 mistral 2 8725
64 openchat/openchat-3.5-0106 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 26858
65 openchat/openchat-3.5-1210 Wang et al. (2024a); OpenAI et al. (2024) mistralai/Mistral-7B-v0.1 mistral 2 2123
@@
72 Nexusflow/NexusRaven-V2-13B Nexusflow.ai team (2023); Rozière et al. (2024) codellama/CodeLlama-13b-Instruct-hf others 2 3966
73 haoranxu/ALMA-13B-Pretrain Xu et al. (2024a, b) meta-llama/Llama-2-13b-hf others 2 5510
@@ 77 Locutusque/Orca-2-13b-SFT_v5 microsoft/Orca-2-13b others 2 2191
80 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) upstage/SOLAR-10.7B-v1.0 others 2 67725
@@
82 deepseek-ai/DeepSeek-Prover-V1.5-RL Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-SFT deepseek 3 12223
@@
86 Intel/neural-chat-7b-v3-3 Yu et al. (2024) Intel/neural-chat-7b-v3-1 mistral 3 166226
@@
Table 7: List of 59 models in the experiments for layer-wise models. “ID” denotes the index sorted in descending order of downloads; “Model Name” denotes the name of the model; “Model Type” denotes the model classification defined in Oyama et al. (2025); “Layers” denotes the number of layers in the model; “DLs” denotes the total number of downloads.
@@
1 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 32 3565248
2 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 2058011
3 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 32 1750089
@@
5 meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023b) llama-2 32 1402244
6 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistral 32 1370245
7 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 32 1294737
@@
10 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 32 675850
@@
12 bigcode/starcoder2-3b Beltagy et al. (2020); Dao et al. (2022); Bavarian et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) others 30 445316
13 lmsys/vicuna-7b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 32 368410
14 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 32 294019
15 meta-llama/Llama-2-13b-chat-hf Touvron et al. (2023b) llama-2 40 266090
16 google/gemma-2b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 18 256798
17 EleutherAI/gpt-neo-1.3B Gao et al. (2020); Black et al. (2021) others 24 243332
18 EleutherAI/gpt-j-6b Gao et al. (2020); Wang and Komatsuzaki (2021); Wang (2021); Su et al. (2023) gptj 28 241435
@@
20 EleutherAI/gpt-neo-2.7B Gao et al. (2020); Black et al. (2021) others 32 205503
@@
22 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 179952
@@ 23 DeepMount00/Llama-3-8b-Ita llama-3 32 179401
24 Intel/neural-chat-7b-v3-3 Yu et al. (2024) mistral 32 166226
25 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 26 157405
@@
27 facebook/opt-1.3b Brown et al. (2020); Zhang et al. (2022) opt 24 139758
28 Qwen/Qwen1.5-1.8B Bai et al. (2023) others 24 137256
29 codellama/CodeLlama-7b-Instruct-hf Rozière et al. (2024) llama-2 32 133523
@@
31 Qwen/Qwen1.5-7B Bai et al. (2023) others 32 122768
32 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 40 120871
33 microsoft/phi-1_5 Li et al. (2023) others 24 112476
34 NousResearch/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 104388
35 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023); Almazrouei et al. (2023); Penedo et al. (2023) falcon 32 104010
36 google/gemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 18 103851
@@ 34 NousResearch/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 32 104388
38 mosaicml/mpt-7b-chat Henry et al. (2020); Press et al. (2022); Dao et al. (2022); MosaicML NLP Team (2023) others 32 88069
39 Qwen/CodeQwen1.5-7B-Chat Bai et al. (2023) others 32 77169
40 GritLM/GritLM-7B Muennighoff et al. (2025) mistral 32 72991
41 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 28 72047
42 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) others 48 67725
43 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 28 66776
44 codellama/CodeLlama-7b-hf Rozière et al. (2024) llama-2 32 66040
45 facebook/opt-2.7b Brown et al. (2020); Zhang et al. (2022) opt 32 61450
46 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) llama-3 32 55718

### v2

@@ 33 DeepMount00/Llama-3-8b-Ita meta-llama/Meta-Llama-3-8B llama-3 2 179401
37 openchat/openchat-3.6-8b-20240522 arxiv:2309.11235 meta-llama/Meta-Llama-3-8B llama-3 2 10464
38 rinna/llama-3-youko-8b gpt-neox-library; rinna-llama-3-youko-8b; llama3modelcard; arxiv:2404.01657 meta-llama/Meta-Llama-3-8B llama-3 2 1468
@@
40 lightblue/suzume-llama-3-8B-multilingual arxiv:2405.12612 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 16442
@@
42 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct waktaversellama3modelcard; llama3modelcard meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2305
43 shenzhi-wang/Llama3-8B-Chinese-Chat shenzhi-wang-2024 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 55718
@@ 44 SJ-Donald/llama3-passthrough-chat meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 2241
45 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA arxiv:2312.09993; llama3modelcard; arxiv:2405.07101 meta-llama/Meta-Llama-3-8B-Instruct llama-3 2 5995
46 NousResearch/Hermes-2-Theta-Llama-3-8B Hermes-2-Theta-Llama-3-8B NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 12392
47 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B arxiv:2404.00495 NousResearch/Hermes-2-Pro-Llama-3-8B llama-3 2 10273
@@
52 cypienai/cymist-2-v02-SFT arxiv:1910.09700 mistralai/Mistral-7B-v0.1 mistral 2 2717
@@
54 HuggingFaceH4/zephyr-7b-alpha arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistralai/Mistral-7B-v0.1 mistral 2 12911
55 HuggingFaceH4/zephyr-7b-beta arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistralai/Mistral-7B-v0.1 mistral 2 294019
56 ibm/merlinite-7b arxiv:2403.01081 mistralai/Mistral-7B-v0.1 mistral 2 11156
@@
58 Intel/neural-chat-7b-v3-1 arxiv:2306.02707 mistralai/Mistral-7B-v0.1 mistral 2 3976
59 kaist-ai/mistral-orpo-capybara-7k arxiv:2403.07691 mistralai/Mistral-7B-v0.1 mistral 2 4821
@@
61 mistralai/Mistral-7B-Instruct-v0.1 arxiv:2310.06825 mistralai/Mistral-7B-v0.1 mistral 2 1370245
62 NousResearch/Hermes-2-Pro-Mistral-7B Hermes-2-Pro-Mistral-7B mistralai/Mistral-7B-v0.1 mistral 2 14586
63 NousResearch/Nous-Hermes-2-Mistral-7B-DPO Nous-Hermes-2-Mistral-7B-DPO mistralai/Mistral-7B-v0.1 mistral 2 8725
64 openchat/openchat-3.5-0106 arxiv:2309.11235; arxiv:2303.08774 mistralai/Mistral-7B-v0.1 mistral 2 26858
65 openchat/openchat-3.5-1210 arxiv:2309.11235; arxiv:2303.08774 mistralai/Mistral-7B-v0.1 mistral 2 2123
@@
72 Nexusflow/NexusRaven-V2-13B nexusraven; arxiv:2308.12950 codellama/CodeLlama-13b-Instruct-hf others 2 3966
73 haoranxu/ALMA-13B-Pretrain arxiv:2309.11674; arxiv:2401.08417 meta-llama/Llama-2-13b-hf others 2 5510
@@ 77 Locutusque/Orca-2-13b-SFT_v5 microsoft/Orca-2-13b others 2 2191
80 upstage/SOLAR-10.7B-Instruct-v1.0 arxiv:2312.15166; arxiv:2403.19270 upstage/SOLAR-10.7B-v1.0 others 2 67725
@@
82 deepseek-ai/DeepSeek-Prover-V1.5-RL arxiv:2408.08152 deepseek-ai/DeepSeek-Prover-V1.5-SFT deepseek 3 12223
@@
86 Intel/neural-chat-7b-v3-3 arxiv:2309.12284 Intel/neural-chat-7b-v3-1 mistral 3 166226
@@
Table 6: List of 50 (+9) models in the experiments for quantization and layer-wise models. “ID” denotes the index sorted in descending order of downloads; “Model Name” denotes the name of the model; “Model Type” denotes the model classification defined in Oyama et al. (2025); “Layers” denotes the number of layers in the model; “DLs” denotes the total number of downloads.
@@
1 mistralai/Mistral-7B-Instruct-v0.2 arxiv:2310.06825 mistral 32 3565248
2 meta-llama/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 32 2058011
3 mistralai/Mistral-7B-v0.1 arxiv:2310.06825 mistral 32 1750089
@@
5 meta-llama/Llama-2-7b-chat-hf arxiv:2307.09288 llama-2 32 1402244
6 mistralai/Mistral-7B-Instruct-v0.1 arxiv:2310.06825 mistral 32 1370245
7 meta-llama/Llama-2-7b-hf arxiv:2307.09288 llama-2 32 1294737
@@
10 meta-llama/Meta-Llama-3-8B llama3modelcard llama-3 32 675850
@@
12 bigcode/starcoder2-3b arxiv:2004.05150; arxiv:2205.14135; arxiv:2207.14255; arxiv:2305.13245; arxiv:2402.19173 others 30 445316
13 lmsys/vicuna-7b-v1.5 arxiv:2307.09288; arxiv:2306.05685 llama-2 32 368410
14 HuggingFaceH4/zephyr-7b-beta arxiv:2305.14233; arxiv:2310.16944; arxiv:2310.01377; arxiv:2305.18290 mistral 32 294019
15 meta-llama/Llama-2-13b-chat-hf arxiv:2307.09288 llama-2 40 266090
16 google/gemma-2b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 18 256798
17 EleutherAI/gpt-neo-1.3B arxiv:2101.00027; gpt-neo others 24 243332
18 EleutherAI/gpt-j-6b arxiv:2101.00027; gpt-j; mesh-transformer-jax; arxiv:2104.09864 gptj 28 241435
@@
20 EleutherAI/gpt-neo-2.7B arxiv:2101.00027; gpt-neo others 32 205503
@@
22 tiiuae/falcon-7b-instruct arxiv:1911.02150; arxiv:2005.14165; arxiv:2205.14135; arxiv:2104.09864; falcon40b; arxiv:2306.01116 falcon 32 179952
@@ 23 DeepMount00/Llama-3-8b-Ita llama-3 32 179401
24 Intel/neural-chat-7b-v3-3 arxiv:2309.12284 mistral 32 166226
25 openlm-research/open_llama_3b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 26 157405
@@
27 facebook/opt-1.3b arxiv:2005.14165; arxiv:2205.01068 opt 24 139758
28 Qwen/Qwen1.5-1.8B arxiv:2309.16609 others 24 137256
29 codellama/CodeLlama-7b-Instruct-hf arxiv:2308.12950 llama-2 32 133523
@@
31 Qwen/Qwen1.5-7B arxiv:2309.16609 others 32 122768
32 meta-llama/Llama-2-13b-hf arxiv:2307.09288 llama-2 40 120871
33 microsoft/phi-1_5 arxiv:2309.05463 others 24 112476
34 NousResearch/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 32 104388
35 tiiuae/falcon-7b arxiv:1911.02150; arxiv:2005.14165; arxiv:2101.00027; arxiv:2205.14135; arxiv:2104.09864; falcon40b; arxiv:2306.01116 falcon 32 104010
36 google/gemma-2b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 18 103851
@@ 34 NousResearch/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 32 104388
38 mosaicml/mpt-7b-chat arxiv:2010.04245; arxiv:2108.12409; arxiv:2205.14135; MosaicML2023Introducing others 32 88069
39 Qwen/CodeQwen1.5-7B-Chat arxiv:2309.16609 others 32 77169
40 GritLM/GritLM-7B arxiv:2402.09906 mistral 32 72991
41 google/gemma-7b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2305.14314; arxiv:2312.11805 gemma 28 72047
42 upstage/SOLAR-10.7B-Instruct-v1.0 arxiv:2312.15166; arxiv:2403.19270 others 48 67725
43 google/gemma-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 28 66776
44 codellama/CodeLlama-7b-hf arxiv:2308.12950 llama-2 32 66040
45 facebook/opt-2.7b arxiv:2005.14165; arxiv:2205.01068 opt 32 61450
46 shenzhi-wang/Llama3-8B-Chinese-Chat shenzhi-wang-2024 llama-3 32 55718